OpenStudy (anonymous):

\[\int\limits_{1/2}^{2}({e^(1/x) \over x^2})dx\]

OpenStudy (anonymous):

\[\int\limits_{1/2}^{2} {e^{1/x} \over x^2}dx\]

OpenStudy (anonymous):

ignore that first bit

OpenStudy (anonymous):

12

OpenStudy (anonymous):

can you tell me how you got that?

OpenStudy (mr.math):

Use a substitution \(u=\frac{1}{x}\).

OpenStudy (anonymous):

ok, this is what I've tried: \[\int\limits_{1/2}^{2}{1/x \over 2 \ln {x}} dx = \int\limits_{-\ln2}^{\ln2}{1 \over 2u}du = {1 \over 2}\ln |u|]_{-\ln2}^{\ln 2} = 0\]

OpenStudy (anonymous):

that last step, btw, was (1/2) (ln|ln 2| - ln|-ln 2|) = (1/2)(ln|ln 2| - ln|ln 2|) = 0

OpenStudy (mr.math):

It seems like you've taken ln of the top and bottom and you're not allowed to do that, it's wrong. Note for example that \(\frac{1}{1}\ne \frac{\ln{1}}{\ln{1}}\). Try to use the substitution I wrote above, it will work.

OpenStudy (anonymous):

ok, I'll give it a go

OpenStudy (anonymous):

@everyone, is the equation editor still functionable? i think it's not,

OpenStudy (anonymous):

oh, so you can take ln of equations but not expressions, is that what it was?

OpenStudy (anonymous):

try to use the equation editor. . i think it's not working. . the tab beside draw tab

OpenStudy (mr.math):

True, you can take ln of both sides of an equation. But you can't take ln of the top and bottom of an expression.

OpenStudy (mr.math):

\[\text{It's working with me.}\]

OpenStudy (anonymous):

no it's not with me. . i don't know what happened. .

OpenStudy (king):

\[Its working!!!\]

OpenStudy (mr.math):

Refresh your page.

OpenStudy (anonymous):

i can't even look unto equations

OpenStudy (anonymous):

I got it :) \[e^2-\sqrt{e}\] And my equation thing's working

OpenStudy (anonymous):

i think it's okay now. . . thank you guys. .

OpenStudy (mr.math):

You've got it right!

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (mr.math):

Anytime.